### v1

@@ 2 Optimal stopping under ambiguity
Suppose that Assumptions 2 and 2 hold. Let $(V_{t}^{x})_{t\in[0,T]}$ be given in (2.2) (see Remark 2.5) and let $(Y_{t}^{x})_{t\in[0,T]}$ be the first component of the unique solution to the reflected BSDE (2.4) with the lower obstacle $(R(X_{t}^{x}))_{t\in[0,T]}$ (see Remark 2.7). Then, $V_{t}^{x}=Y_{t}^{x}$, $\mathbb{P}$-a.s. for all $t\in[0,T]$. In particular, the stopping time $\tau_{t}^{*,x}\in{\mathcal{}T}_{t}$, defined by
(2.5) $\tau_{t}^{*,x}:=\inf\{s\geq t\,|\,Y_{t}^{x}\leq R(X_{t}^{x})\}\wedge T,$
is optimal to the robust stopping problem $V^{x}$.
The penalization method is a standard approach for establishing the existence of solutions to reflected BSDEs (see, e.g., [21, 39, 54]). We introduce a sequence of penalized BSDEs and remark on the convergence of their solutions to that of the reflected BSDE given (2.4).
To that end, set for every $N\in\mathbb{N}$ and $(\omega,t,y,z)\in\Omega\times[0,T]\times\mathbb{R}\times\mathbb{R}^{d}$ by
(2.6) $F_{t}^{x;N}(\omega,y,z):=F_{t}^{x}(\omega,y,z)+N(R\big(X_{t}^{x}(\omega)\big)-y)^{+},$
where $F^{x}$ is given in (2.3) and $(a)^{+}:=\max\{a,0\}$ for $a\in\mathbb{R}$. Then we denote for every $N\in\mathbb{N}$ by $(Y_{t}^{x;N},Z_{t}^{x;N})_{t\in[0,T]}$ a couple of processes satisfying that
(2.7) $Y_{t}^{x;N}=R(X_{T}^{x})+\int_{t}^{T}F_{s}^{x;N}(Y_{s}^{x;N},Z_{s}^{x;N})ds-\int_{t}^{T}Z_{s}^{x;N}dB_{s},\;\;\mbox{for $t\in[0,T]$}.$
@@
Under Assumptions 2 and 2, the parameters of the BSDE (2.7) satisfy all the conditions given in [49, Section 3]. Hence, we recognize:
@@
    For every $N\in\mathbb{N}$ there exists a unique solution $(Y^{x;N}_{t},Z_{t}^{x;N})_{t\in[0,T]}\in\mathbb{S}^{2}(\mathbb{R})\times\mathbb{L}^{2}(\mathbb{R}^{d})$ of the BSDE (2.7) (see [49, Theorem 3.1]).
@@
    Moreover, if we set $K_{t}^{x;N}:=N\int_{0}^{t}(R(X_{s}^{x})-Y_{s}^{x;N})^{+}ds$ for $t\in[0,T]$, then it follows from [20, Section 6., Eq. (16)] that there exists some constant $C>0$ such that for every $N\in\mathbb{N}$, $\|Y^{x;N}\|_{\mathbb{S}^{2}}^{2}+\|Z^{x;N}\|_{\mathbb{L}^{2}}^{2}+\|K_{T}^{x;N}\|_{L^{2}}^{2}\leq C.$
@@
    Lastly, we recall that $(Y_{t}^{x},Z_{t}^{x},K_{t}^{x})_{t\in[0,T]}$ is the unique solution to the reflected $g$-BSDE (2.4) (see Remark 2.7). Then, it follows from [39, Lemma 3.2 & Theorem 3.3] that [^1] We say $Z\in\mathbb{L}^{2}(\mathbb{R}^{d})$ is the weak limit of $(Z^{n})_{n\in\mathbb{N}}\subseteq\mathbb{L}^{2}(\mathbb{R}^{d})$ if for every $\phi\in\mathbb{L}^{2}(\mathbb{R}^{d})$, it holds that $\langle Z^{n},\phi\rangle_{\mathbb{P}\otimes dt}\to\langle Z,\phi\rangle_{\mathbb{P}\otimes dt}$ as $n\to\infty$, where the inner product is defined by $\langle L,M\rangle_{\mathbb{P}\otimes dt}:=\mathbb{E}[\int_{0}^{T}\langle L_{t},M_{t}\rangle dt]$ for $L,M\in\mathbb{L}^{2}(\mathbb{R}^{d})$. Similarly, the weak limit in $L^{2}({\mathcal{}F}_{t};\mathbb{R}^{d})$ is defined w.r.t. the inner product $\langle\xi,\eta\rangle_{\mathbb{P}}:=\mathbb{E}[\langle\xi,\eta\rangle]$ for $\xi,\eta\in L^{2}({\mathcal{}F}_{t};\mathbb{R}^{d})$. $Y^{x}$ is the strong limit of $(Y^{x;N})_{N\in\mathbb{N}}$ in $\mathbb{L}^{2}(\mathbb{R})$ (i.e., as $N\to\infty$ $\|Y^{x;N}-Y^{x}\|_{\mathbb{L}^{2}}\to 0$), $Z^{x}$ is the weak limit of $(Z^{x;N})_{N\in\mathbb{N}}$ in $\mathbb{L}^{2}(\mathbb{R}^{d})$, and for each $t\in[0,T]$ $K_{t}^{x}$ is the weak limit of $K_{t}^{x;N}$ in ${L}^{2}({\mathcal{}F}_{t};\mathbb{R})$.
The following proposition shows that for each $N\in\mathbb{N}$ the solution to the penalized BSDE (2.7) can be represented by a certain optimal stochastic control problem under ambiguity. The corresponding proof is presented in Section 6.1.
@@
Suppose that Assumptions 2 and 2 hold. Let $N\in\mathbb{N}$ be given. Denote by $Y^{x;N}$ the first component of the unique solution to (2.7). Then $Y^{x;N}$ admits a representation of the robust control optimization problem in the following sense: Let $\mathcal{}A$} be the set of all $\mathbb{F}$-progressively measurable processes $\alpha=(\alpha_{t})_{t\in[0,T]}$ with values in $\{0,1\}$. Set for every $t\in[0,T]$ and $N\in\mathbb{N}$
@@ 3 Exploratory framework: approximation of optimal stopping under ambiguity
Based on the results in Section 2, we are able to show that for sufficiently large $N\in\mathbb{N}$, the optimal stopping problem $V^{x}(=Y^{x})$ under ambiguity in (2.2) (see also Proposition 2.8) can be approximated by the optimal stochastic control problem $Y^{x;N}$ under ambiguity (see Proposition 2.10). The proofs of all the results in this section are presented in Section 6.2.
We introduce an exploratory framework of [66, 67] into $Y^{x;N}$. In particular, we aim to study a robust analogue of the optimal exploratory stopping framework in [15]. To that end, let $\Pi$ be the set of all $\mathbb{F}$-progressively measurable processes $\pi=(\pi_{t})_{t\in[0,T]}$ taking values in $[0,1]$, i.e., an exploratory version of the $\{0,1\}$-valued controls set $\mathcal{}A$} appearing in Proposition 2.10.
Then let ${\mathcal{}H}:[0,1]\ni a\to{\mathcal{}H}(a)\in\mathbb{R}$ be the binary differential entropy defined by
(3.1) ${\mathcal{}H}(a):=a\log(a)+(1-a)\log(1-a)\quad\mbox{for $a\in(0,1)$},$
with the convention that ${\mathcal{}H}(0):=\lim_{a\downarrow 0}{\mathcal{}H}(a)=0$ and ${\mathcal{}H}(1):=\lim_{a\uparrow 1}{\mathcal{}H}(a)=0$.
@@
We can then describe the decision-maker’s optimal exploratory control problem $\overline{V}^{x;N,\lambda}:=(\overline{V}_{t}^{x;N,\lambda})_{t\in[0,T]}$ under ambiguity for any $N\in\mathbb{N}$ and $\lambda>0$:
@@
$\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi}:=$ $\int_{t}^{T}e^{-\int_{t}^{s}(\beta_{u}+N\pi_{u})du}\big(r(X_{s}^{x})+R(X_{s}^{x})\,N\pi_{s}-\lambda{\mathcal{}H}(\pi_{s})\big)$
@@
We note that the differential entropy $\mathcal{}H$} given in (3.1) is strictly convex and bounded on $[0,1]$. Moreover, since all the exploratory control $\pi\in\Pi$ is uniformly bounded by $[0,1]$, by using the same arguments presented for Remark 2.5, we have that $\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi}\in L^{2}({\mathcal{}F}_{T};\mathbb{R})$ for all $N\in\mathbb{N}$, $\lambda>0$, and $\pi\in\Pi$. Therefore, $\overline{V}^{x;N,\lambda}$ given in (3.2) is well-defined for all $N\in\mathbb{N}$ and $\lambda>0$.
@@
Assume that the probability space $(\Omega,{\mathcal{}F},\mathbb{P})$ supports a uniformly distributed random variable $U$ with values in $[0,1]$ which is independent of the fixed Brownian motion $B$. Then we are able to see that each exploratory control $\pi\in\Pi$ generates a Bernoulli-distributed (randomized) process under drift ambiguity. Indeed, we recall the variational characterization of $g$-expectation in Remark 2.2 with the map $\hat{g}:\Omega\times[0,T]\times\mathbb{R}^{d}\to\mathbb{R}$ and the set ${\mathcal{}B}^{g}$. Then, for all $N\in\mathbb{N}$, $\lambda>0$, and $t\in[0,T]$, we can rewrite the conditional $g$-expectation value ${\mathcal{}E}_{t}^{g}[\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi}]$ given in (3.2) as the following strong formulation for drift ambiguity under $\mathbb{P}$ (see [1, Section 5]):
(3.3) ${\mathcal{}E}_{t}^{g}[\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi}]=\operatorname*{ess\,inf}_{\vartheta\in{\mathcal{}B}^{g}}\mathbb{E}_{t}[\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi,\vartheta}+\int_{t}^{T}\hat{g}(s,\vartheta_{s})ds],$
@@
Then by using the random variable $U$ and its independence with the filtration $\mathbb{F}$ generated by $B$, we can apply the Blackwell–Dubins lemma (see [8]) to ensure that there exists a (randomized) process $(\widetilde{\alpha}_{t})_{t\in[0,T]}$ such that for every $t\in[0,T]$, $\mathbb{P}$-a.s.,
@@
In order to characterize $\overline{V}^{x;N,\lambda}$ given in (3.2), we first collect several preliminary results concerning the following auxiliary BSDE formulations: Recall that $F^{x}$ is given in (2.3). Set for every $\pi\in\Pi$ and $(\omega,t,y,z)\in\Omega\times[0,T]\times\mathbb{R}\times\mathbb{R}^{d}$
@@
Then, consider the (controlled) processes $(\overline{Y}_{t}^{x;N,\lambda,\pi},\overline{Z}_{t}^{x;N,\lambda,\pi})_{t\in[0,T]}$ satisfying
@@
    Since $(\pi_{t})_{t\in[0,T]}\in\Pi$ and $({\mathcal{}H}(\pi_{t}))_{t\in[0,T]}$ are uniformly bounded (see Remark 3.1), we are able to see that the parameters of (3.5) satisfy all the conditions given in [49, Section 3]. Therefore, there exists a unique solution $(\overline{Y}_{t}^{x;N,\lambda,\pi},\overline{Z}_{t}^{x;N,\lambda,\pi})_{t\in[0,T]}\in\mathbb{S}^{2}(\mathbb{R})\times\mathbb{L}^{2}(\mathbb{R}^{d})$ to (3.5).
@@
Moreover, set for every $N\in\mathbb{N}$, $\lambda>0$, and $(\omega,t,y,z)\in\Omega\times[0,T]\times\mathbb{R}\times\mathbb{R}^{d}$ by
@@
Then consider the couple of processes $(\overline{Y}_{t}^{x;N,\lambda},\overline{Z}_{t}^{x;N,\lambda})_{t\in[0,T]}$ satisfying
@@
In the following theorem, the optimal exploratory control problem $\overline{V}^{x;N,\lambda}$ under ambiguity and its optimal control are characterized via the auxiliary BSDE given in (3.8).
@@
    Moreover, recall $\overline{V}^{x;N,\lambda}$ is given in (3.2). Then it holds for every $t\in[0,T]$ that $\overline{Y}_{t}^{x;N,\lambda}=\overline{V}_{t}^{x;N,\lambda}={\mathcal{}E}_{t}^{g}[\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi^{*,x;N,\lambda}}]$ $\mathbb{P}$-a.s., where the optimizer $\pi^{*,x;N,\lambda}:=(\pi^{*,x;N,\lambda}_{t})_{t\in[0,T]}\in\Pi$ is given by
@@
The following theorem is devoted to showing the comparison and stability results between the exploratory and non-exploratory optimal control problems characterized in Proposition 2.10 and Theorem 3.4.
@@
As a consequence of Theorem 3.5, the following corollary establishes the asymptotic behavior of the optimal exploratory control derived in Theorem 3.4 into the optimal non-exploratory control derived in Proposition 2.10.
Corollary 3.6.
Suppose that Assumptions 2 and 2 hold. For each $N\in\mathbb{N}$ and $\lambda>0$, let $\alpha^{*,x;N}\in{\mathcal{}A}$ and $\pi^{*,x;N,\lambda}\in\Pi$ be defined as in (2.8) and (3.9), respectively. Then it holds that for every $N\in\mathbb{N}$,
(3.12) $\big\|\alpha^{*,x;N}-\pi^{*,x;N,\lambda}\big\|_{\mathbb{L}^{1}}\to 0\quad\mbox{as $\lambda\downarrow 0$},$
i.e., for any $N\in\mathbb{N}$, $\pi^{*,x;N,\lambda}$ strongly converges to $\alpha^{*,x;N}$ in the set of all $\mathbb{F}$ progressively measurable processes endowed with the norm $\|\cdot\|_{\mathbb{L}^{1}}$, as $\lambda\downarrow 0$.

### v2

@@ 2 Optimal stopping under ambiguity
Suppose that Assumptions 2 and 2 hold. Let $V^{x}$ be given in (2.2) (see Remark 2.5) and let $Y^{x}$ be the first component of the unique solution to the reflected BSDE (2.4) with the obstacle $(R(X_{t}^{x}))_{t\in[0,T]}$ (see Remark 2.7). Then, we have that $V_{t}^{x}=Y_{t}^{x}$ $\mathbb{P}$-a.s. for all $t\in[0,T]$. In particular, the stopping time
(2.5) $\tau_{t}^{*,x}:=\inf\{s\geq t:Y_{s}^{x}\leq R(X_{s}^{x})\}\wedge T\in{\mathcal{}T}_{t}$
is optimal to $V_{t}^{x}$, i.e., ${\mathcal{}E}^{g}_{t}[\operatorname{I}_{t}^{x;\tau^{*,}_{t}}]=V_{t}^{x}$ $\mathbb{P}$-a.s..
In the remainder of this section, we introduce auxiliary robust stochastic control problems arising from the penalization method which is used to establish the existence of solutions to reflected BSDEs (see, e.g., [el1997backward, lepeltier2005penalization, peng2005smallest]).
To that end, define for every $N\in\mathbb{N}$ and $(\omega,t,y,z)\in\Omega\times[0,T]\times\mathbb{R}\times\mathbb{R}^{d}$ by
(2.6) $F_{t}^{x;N}(\omega,y,z):=F_{t}^{x}(\omega,y,z)+N\big(R(X_{t}^{x}(\omega))-y\big)^{+},$
where $F^{x}$ is given in (2.3) and $(a)^{+}:=\max\{a,0\}$ for $a\in\mathbb{R}$. Then we denote for every $N\in\mathbb{N}$ by $(Y^{x;N},Z^{x;N})$ a couple of processes satisfying
(2.7) $Y_{t}^{x;N}=R(X_{T}^{x})+\int_{t}^{T}F_{s}^{x;N}(Y_{s}^{x;N},Z_{s}^{x;N})ds-\int_{t}^{T}Z_{s}^{x;N}dB_{s},\quad\mbox{$t\in[0,T]$}.$
@@
Under Assumptions 2 and 2, the parameters of the BSDE (2.7) satisfy all the conditions given in [pardoux1990adapted, Section 3]. Hence, the following hold:
@@
    For every $N\in\mathbb{N}$, there exists a unique solution $(Y^{x;N},Z^{x;N})\in\mathbb{S}^{2}(\mathbb{R})\times\mathbb{L}^{2}(\mathbb{R}^{d})$ of the BSDE (2.7); see [pardoux1990adapted, Theorem 3.1].
@@
    If we set $K_{t}^{x;N}:=N\int_{0}^{t}(R(X_{s}^{x})-Y_{s}^{x;N})^{+}ds$ for $t\in[0,T]$, then it follows from [el1997reflected, Section 6., Eq. (16)] that there exists some constant $C>0$ such that for every $N\in\mathbb{N}$, $\|Y^{x;N}\|_{\mathbb{S}^{2}}^{2}+\|Z^{x;N}\|_{\mathbb{L}^{2}}^{2}+\|K_{T}^{x;N}\|_{L^{2}}^{2}\leq C.$
@@
    We recall that $(Y^{x},Z^{x},K^{x})$ is the unique solution to (2.4); see Remark 2.7. We have by [el1997reflected, Section 6] that as $N\to\infty$,
    $\|Y^{x;N}-Y^{x}\|_{\mathbb{S}^{2}}+\|Z^{x;N}-Z^{x}\|_{\mathbb{L}^{2}}+\|K^{x;N}-K^{x}\|_{\mathbb{S}^{2}}\to 0.$
    In particular, it holds that $Y_{t}^{x;N}\uparrow Y_{t}^{x}$ $\mathbb{P}$-a.s. for all $t\in[0,T]$
The following proposition shows that for each $N\in\mathbb{N}$ the solution to the penalized BSDE (2.7) can be represented by a robust stochastic control problem under ambiguity. The corresponding proof is presented in Section 6.1.
@@
Suppose that Assumptions 2 and 2 hold. Let $N\in\mathbb{N}$ be given. Denote by $Y^{x;N}$ the first component of the unique solution to (2.7). Then $Y^{x;N}$ admits a representation of the robust stochastic control problem in the following sense: Let $\mathcal{}A$} be the set of all $\mathbb{F}$-progressively measurable processes $\alpha=(\alpha_{t})_{t\in[0,T]}$ with values in $\{0,1\}$. Set for every $t\in[0,T]$ and $N\in\mathbb{N}$
@@ 3 Exploratory framework: approximation of optimal stopping under ambiguity
In the previous section, we show that for sufficiently large $N\in\mathbb{N}$, the optimal stopping problem $V^{x}(=Y^{x})$ under ambiguity in (2.2) can be approximated by the robust stochastic control problem $Y^{x;N}$ under ambiguity.
In this section, we introduce an exploratory version of $Y^{x;N}$, following [wang2020reinforcement, wang2020continuous]. In particular, we aim to study a robust analogue of the optimal exploratory stopping framework in [dai2024learning]. To that end, let $\Pi$ be the set of all $\mathbb{F}$-progressively measurable processes $\pi=(\pi_{t})_{t\in[0,T]}$ taking values in $[0,1]$, i.e., an exploratory version of the $\{0,1\}$-valued controls set $\mathcal{}A$} appearing in Proposition 2.10.
Then we define ${\mathcal{}H}:[0,1]\ni a\mapsto{\mathcal{}H}(a)\in\mathbb{R}$ by
(3.1) ${\mathcal{}H}(a):=a\log(a)+(1-a)\log(1-a),$
with the convention that ${\mathcal{}H}(0):=\lim_{a\downarrow 0}{\mathcal{}H}(a)=0$ and ${\mathcal{}H}(1):=\lim_{a\uparrow 1}{\mathcal{}H}(a)=0$. Typically, ${\mathcal{}H}(a)$ is called the binary differential entropy.
@@
We can then describe the robust exploratory control problem under ambiguity: for any $N\in\mathbb{N}$ and $\lambda>0$:
@@
$\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi}:=$ $\int_{t}^{T}e^{-\int_{t}^{s}(\beta_{u}+N\pi_{u})du}\big(r(X_{s}^{x})+R(X_{s}^{x})\,N\pi_{s}-\lambda{\mathcal{}H}(\pi_{s})\big)ds$
@@
We note that $\mathcal{}H$} in (3.1) is strictly convex and bounded on $[0,1]$. Moreover, since every $\pi\in\Pi$ is uniformly bounded by $[0,1]$, by the same arguments presented for Remark 2.5, we have that $\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi}\in L^{2}({\mathcal{}F}_{T};\mathbb{R})$ for all $N\in\mathbb{N}$, $\lambda>0$, and $\pi\in\Pi$. Therefore, $\overline{V}^{x;N,\lambda}$ given in (3.2) is well-defined for all $N\in\mathbb{N}$ and $\lambda>0$.
@@
Assume that the probability space $(\Omega,{\mathcal{}F},\mathbb{P})$ supports a uniformly distributed random variable $U$ with values in $[0,1]$ which is independent of the fixed Brownian motion $B$. Then we are able to see that every exploratory control $\pi\in\Pi$ generates a Bernoulli-distributed (randomized) process under drift ambiguity. Indeed, we recall the variational characterization of $g$-expectation in Remark 2.2 with the map $\hat{g}:\Omega\times[0,T]\times\mathbb{R}^{d}\to\mathbb{R}$ and the set ${\mathcal{}B}^{g}$. Then, for all $N\in\mathbb{N}$, $\lambda>0$, and $t\in[0,T]$, we can rewrite the conditional $g$-expectation value ${\mathcal{}E}_{t}^{g}[\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi}]$ given in (3.2) as the following strong formulation for drift ambiguity under $\mathbb{P}$ (see [bartl2024numerical, Section 5]):
(3.3) ${\mathcal{}E}_{t}^{g}[\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi}]=\operatorname*{ess\,inf}_{\vartheta\in{\mathcal{}B}^{g}}\mathbb{E}_{t}\bigg[\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi,\vartheta}+\int_{t}^{T}\hat{g}(s,\vartheta_{s})ds\bigg],$
@@
Then by using the random variable $U$ and its independence with the filtration $\mathbb{F}$ generated by $B$, we can apply the Blackwell–Dubins lemma (see [blackwell1983extension]) to ensure that there exists a randomized process $(\widetilde{\alpha}_{t})_{t\in[0,T]}$ such that for every $t\in[0,T]$, $\mathbb{P}$-a.s.,
@@
In order to characterize $\overline{V}^{x;N,\lambda}$ given in (3.2), we first collect several preliminary results concerning the following auxiliary BSDE formulations: Recall that $F^{x}$ is given in (2.3). Denote for every $\pi\in\Pi$ and $(\omega,t,y,z)\in\Omega\times[0,T]\times\mathbb{R}\times\mathbb{R}^{d}$ by
@@
Then, consider the controlled processes $(\overline{Y}^{x;N,\lambda,\pi},\overline{Z}^{x;N,\lambda,\pi})$ satisfying
@@
    Since $(\pi_{t})_{t\in[0,T]}\in\Pi$ and $({\mathcal{}H}(\pi_{t}))_{t\in[0,T]}$ are uniformly bounded (see Remark 3.1), we are able to see that the parameters of (3.5) satisfy all the conditions given in [pardoux1990adapted, Section 3]. Therefore, there exists a unique solution $(\overline{Y}^{x;N,\lambda,\pi},\overline{Z}^{x;N,\lambda,\pi})\in\mathbb{S}^{2}(\mathbb{R})\times\mathbb{L}^{2}(\mathbb{R}^{d})$ to (3.5).
@@
Moreover, define for every $N\in\mathbb{N}$, $\lambda>0$, and $(\omega,t,y,z)\in\Omega\times[0,T]\times\mathbb{R}\times\mathbb{R}^{d}$ by
@@
Then consider the couple of processes $(\overline{Y}^{x;N,\lambda},\overline{Z}^{x;N,\lambda})$ satisfying
@@
The following theorem shows that the robust exploratory control problem $\overline{V}^{x;N,\lambda}$ is characterized by the BSDE (3.8). The proof is presented in Section 6.2.
@@
    Moreover, recall $\overline{V}^{x;N,\lambda}$ is given in (3.2). Then it holds for every $t\in[0,T]$ that $\overline{Y}_{t}^{x;N,\lambda}=\overline{V}_{t}^{x;N,\lambda}={\mathcal{}E}_{t}^{g}[\overline{\operatorname{J}}_{t}^{x;N,\lambda,\pi^{*,x;N,\lambda}}]$ $\mathbb{P}$-a.s., where $\pi^{*,x;N,\lambda}\in\Pi$ is given by
@@
The following theorem establishes the comparison and stability results between the exploratory and non-exploratory robust control problems characterized in Proposition 2.10 and Theorem 3.4. The proof is presented in Section 6.2.
@@
Having completed the characterization of the robust exploratory control problem, we now introduce a stopping time ${\tau}^{x;N,\lambda}_{t}\in{\mathcal{}T}_{t}$ defined, for every $N\in\mathbb{N}$ and $\lambda>0$, by
(3.12) $\begin{aligned} {\tau}^{x;N,\lambda}_{t}:=&\inf\bigg\{s\geq t:\pi^{*,x;N,\lambda}_{s}\geq\frac{1}{1+2^{N(T-s)}}\bigg\}\wedge T\\ =&\inf\{s\geq t:\overline{Y}_{s}^{x;N,\lambda}-\lambda(T-s)\log 2\leq R(X_{s}^{x})\}\wedge T\quad t\in[0,T],\end{aligned}$
where the equality holds by definition of the standard logistic function $\operatorname{logit}(\cdot)$. We refer to it as a robust exploratory stopping time.
The following theorem shows that the robust exploratory stopping time ${\tau}^{x;N,\lambda}_{t}$ in (3.12) converges to the optimal stopping time $\tau_{t}^{*,x}$ in (2.5) along any sequences $(N_{n},\lambda_{n})$ with $N_{n}\uparrow\infty$ and $\lambda_{n}\downarrow 0$. Moreover, it establishes that the corresponding $g$-expectation value under the robust exploratory stopping time converges to the optimal stopping value $V_{t}^{x}(=Y_{t}^{x})$ in (2.2). The proof can be found in Section 6.2.
Theorem 3.6.
Suppose that Assumptions 2 and 2 hold. Let $(N_{n})_{n\in\mathbb{N}}\subset\mathbb{N}$ and $(\lambda_{n})_{n\in\mathbb{N}}\subset(0,\infty)$ be sequences satisfying $N_{n}\uparrow\infty$ and $\lambda_{n}\downarrow 0$ as $n\to\infty$. Then it holds for every $t\in[0,T]$ that as $n\to\infty$,
$\tau_{t}^{x;N_{n},\lambda_{n}}\to\tau_{t}^{*,x}\quad\mbox{$\mathbb{P}$-a.s.},\quad\mbox{and}\quad\big\|{\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau_{t}^{x;N_{n},\lambda_{n}}}]-V_{t}^{x}\big\|_{L^{2}}\to 0.$
Remark 3.7.
We note that the exploration benefit $\overline{Y}^{x;N,\lambda}-Y^{x;N}$ compensates for the penalization error $Y^{x}-Y^{x;N}$ (see (3.10) and Remark 2.9 (iii)), which does not lead to a direct comparison between $\overline{Y}^{x;N,\lambda}$ and $Y^{x}$. Consequently, it is not immediate whether a stopping time defined, for every $N\in\mathbb{N}$ and $\lambda>0$, by
$\overline{\tau}^{x;N,\lambda}_{t}:=$ $\inf\bigg\{s\geq t:\pi^{*,x;N,\lambda}_{s}\geq\frac{1}{2}\bigg\}\wedge T=\inf\{s\geq t:\overline{Y}_{s}^{x;N,\lambda}\leq R(X_{s}^{x})\}\wedge T\in{\mathcal{}T}_{t},$
converges to $\tau^{*;x}_{t}$ in (2.5), even along sequences $(N_{n},\lambda_{n})$ with $N_{n}\uparrow\infty$ and $\lambda_{n}\downarrow 0$.
To construct a stopping time that converges to $\tau^{*,x}_{t}$, we first establish the bound $\overline{Y}_{t}^{x;N,\lambda}-Y_{t}^{x}\leq\lambda(T-t)\log 2$. This allows us to consider the value $\overline{Y}_{t}^{x;N,\lambda}-\lambda(T-t)\log 2$ which is dominated by $Y^{x}$ and also converges to $Y^{x}$ along $(N_{n},\lambda_{n})$ with $N_{n}\uparrow\infty$ and $\lambda_{n}\downarrow 0$. Then, by establishing the convergence of an increasing sequence of hitting times of continuous processes (see Lemma 6.5), we obtain in Theorem 3.6 that the robust exploratory stopping time $\tau_{t}^{x;N,\lambda}$ in (3.12) converges to $\tau^{*,x}_{t}$. This result also applies to the non-robust (classical) exploratory stopping framework of [dai2024learning].